# CacheFlow: Fast Human Motion Prediction by Cached Normalizing Flow

### v1

@@ Abstract
Many density estimation techniques for 3D human motion prediction require a significant amount of inference time, often exceeding the duration of the predicted time horizon. To address the need for faster density estimation for 3D human motion prediction, we introduce a novel flow-based method for human motion prediction called CacheFlow. Unlike previous conditional generative models that suffer from time efficiency, CacheFlow takes advantage of an unconditional flow-based generative model that transforms a Gaussian mixture into the density of future motions. The results of the computation of the flow-based generative model can be precomputed and cached. Then, for conditional prediction, we seek a mapping from historical trajectories to samples in the Gaussian mixture. This mapping can be done by a much more lightweight model, thus saving significant computation overhead compared to a typical conditional flow model. In such a two-stage fashion and by caching results from the slow flow model computation, we build our CacheFlow without loss of prediction accuracy and model expressiveness. This inference process is completed in approximately one millisecond, making it 4$\times$ faster than previous VAE methods and 30$\times$ faster than previous diffusion-based methods on standard benchmarks such as Human3.6M and AMASS datasets. Furthermore, our method demonstrates improved density estimation accuracy and comparable prediction accuracy to a SOTA method on Human3.6M. Our code and models will be publicly available.
@@ 1 Introduction
The task of 3D human motion prediction is to forecast the future 3D pose sequence given an observed past sequence. Traditional motion prediction methods are often based on deterministic models and can struggle to capture the inherent uncertainty in human movement. Recently, stochastic approaches have addressed this limitation. Stochastic approaches allow models to sample multiple possible future motions. Stochastic human motion prediction methods utilize conditional generative models such as generative adversarial networks (GANs) [18], variational autoencoders (VAEs) [27], and denoising diffusion probabilistic model [24]. However, many stochastic approaches cannot explicitly model the probability density distribution.
Conversely, density estimate-based approaches explicitly model the probability density distribution. In safety-critical applications such as autonomous driving [51] and human-robot interaction [29, 31, 9], a density estimate can represent all possible future motions (not just a few samples) by tracking the volume of density. It can be used to derive guarantees on safety [43, 58, 64].
However, previous density estimation suffers from high computational cost. The expensive computational cost can prohibit applications to real-time use-cases, especially with high dimensional data such as human motions. For instance, kernel density estimation (KDE) [56, 52] requires an exponentially growing number of samples for accurate estimation. Concretely, more than one trillion samples are required for accurate KDE over a 48-dim pose over 100 frames of human motion prediction [60].
In contrast to traditional KDE, recent parametric density estimation approaches use conditional flow-based generative models, including normalizing flows [55, 62, 63] and continuous normalizing flows [13]. These flow-based generative models (“flow-based model” for brevity) directly estimate the density to avoid time-consuming sampling required in KDE. However, inferring the exact probability of possible future motions remains computationally expensive. This is because capturing the full shape of the distribution requires evaluating the probabilities of many potential future motions.
To address this computational limitation, we propose a fast density estimation method based on a flow-based model called "CacheFlow". Our CacheFlow utilizes an unconditional flow-based model for prediction, as illustrated in Figure 1. Since the unconditional flow-based model is independent of past observed motions, its calculation can be precomputed and skipped at inference. This precomputation omits a large portion of computational cost. To achieve further acceleration, our unconditional flow-based model represents transformation between a lightweight conditional base density and the density of future motions. At inference, the density of future motion is estimated by computing the lightweight conditional base density and combining it with the precomputed results of the flow-based model. The inference of our method is approximately one millisecond.
CacheFlow demonstrates comparable accuracy to previous methods on standard stochastic human motion prediction benchmarks, Human3.6M [25] and AMASS [42]. Furthermore, our method estimates density more accurately than previous stochastic human motion prediction methods with KDE. CacheFlow shows improved computational efficiency, making it well-suited for real-time applications. The contributions of this paper are four-fold as follows:
@@ 2.1 Human Motion Prediction
Deterministic approaches. Early approaches on human motion prediction [1, 7, 17, 8, 21, 26, 33] focused on deterministic settings. They predict the most likely motion sequence based on the past motion. A wide range of architectures were proposed including multi-layer perceptron [21], recurrent neural networks [17, 26, 46, 20, 53, 37], convolutional neural networks [33, 50], transformers [1, 10, 48], and graph neural networks (GNNs) [44, 34, 14, 35]. GNN can account for the explicit tree expression of the human skeleton, while other architectures implicitly learn the dependencies between joints.
Stochastic approaches. To capture the inherent uncertainty in human movements, recent works have focused on stochastic human motion prediction to predict multiple likely future motions. The main stream of stochastic methods use generative models for the purpose, such as generative adversarial networks (GANs) [5, 30], variational autoencoder (VAE) [65, 70, 45, 11], and denoising diffusion probabilistic model (DDPM) [4, 12, 66, 61]. To improve the diversity of predictions, diversity-promoting loss [45, 4] or explicit sampling techniques [66] were proposed. In contrast to generative models, anchor-based methods [69, 68] learn a fixed number of anchors corresponding to each prediction to ensure diversity. However, most stochastic methods cannot describe the density of future motions explicitly. This prevents exhaustive or maximum likelihood sampling for practical applications. On the contrary, our method allows for explicit density estimation using normalizing flows [28].
@@ 2.2 Density Estimation
Non-parametric Approach. The representative non-parametric density estimation is kernel density estimation (KDE) [56, 52]. KDE can estimate density by using samples from generative models. However, KDE requires a large number of samples for accurate estimation. Therefore, it often cannot run in real-time.
Parametric Approach. As a representative parametric model, Gaussian mixture models (GMMs) parametrize density with several Gaussian distributions and their mixture weights. Its nature of mixing Gaussian priors limits its ability to generalize to complex data distribution. Another parametric approach with more expressivity is flow-based generative models [28]. By a learned bijective process, normalizing flows (NFs) [55, 62, 63] transform a simple density like the standard normal distribution into a complex data density. Recently, continuous normalizing flows (CNFs) [13, 19] achieve more expressive density than standard normalizing flows via an ODE-based bijective process. While training of CNFs is inefficient due to the optimization of ODE solutions, an efficient training strategy named flow matching [36] was proposed. FlowChain [40] was proposed for fast and efficient density estimation in human trajectory forecasting. FlowChain improves the inference time efficiency by reusing results from the conditional flow-based method while the past sequences are similar. However, with significantly different past sequences, FlowChain’s efficiency can’t hold anymore. Unlike FlowChain, our method can perform fast and efficient inference regardless of past sequences.
@@ 3.2 Normalizing Flow
Normalizing flow [55, 62, 63] is a generative model with explicit density estimation. It follows a bijective mapping $f_{\theta}$ with learnable parameters $\theta$. It transforms a simple base density $q(\bm{z})$ such as a Gaussian distribution into the complex data density $p(\bm{x})$. We can analytically estimate the exact probability via the change-of-variables formula as follows:
@@
where $\mathcal{J}_{f_{\theta}}(z)=\frac{\partial f_{\theta}}{\partial\bm{z}}$ is the Jacobian of $f_{\theta}$ at $\bm{z}$. The parameters $\theta$ of $f_{\theta}$ can be learned by maximizing the likelihood (or conditional likelihood) of samples $\hat{\bm{x}}$ from datasets or minimizing the negative log-likelihood as $\mathcal{L}_{\text{NLL}}=-\log p(\hat{\bm{x}})$. When $\bm{x}$ and $\bm{z}$ are latent codes, normalizing flow is transformed into latent normalizing flow. We follow this pattern in our method. We encode the past human motion into $\bm{x}$ by an ecoder network $\mathcal{E}$ and decode it by a decoder network $\mathcal{D}$:
@@ 3.3 Continuous Normalizing Flow (CNF)
Continuous normalizing flow (CNF) [13, 19] is a normalizing flow variant based on an ordinary differential equation (ODE). CNF defines $t$-continuous path $\bm{z}_{t}$ between the base density space $\bm{z}_{0}\sim q(\bm{z})$ and the data space $\bm{z}_{1}=\bm{x}\sim p(\bm{x})$. This $\bm{z}_{t}$ is defined by the parameterized vector field $\frac{d\bm{z}_{t}}{dt}=v_{\theta}(\bm{z}_{t})$. The data $\bm{x}=\bm{z}_{1}$ is generated via numerical integration of vector field $v_{\theta}(\bm{z}_{t})$ as follows:
[…]
The CNF transformation Equation 4 is denoted as $\bm{x}=f_{\theta}(\bm{z})$ for brevity. Although CNF can be trained by minimizing negative log-likelihood, it is time-consuming due to the numerical integration of ODE.
@@ 3.4 Flow Matching
In order to train the parameterized vector field efficiently, one can leverage the flow matching [36] strategy. As a new training strategy, Flow Matching avoids the numerical integration of ODE by directly optimizing the vector field $v_{\theta}(\bm{z}_{t})$. The objective of flow matching is to match the parameterized vector field $v_{\theta}(\bm{z}_{t})$ to the ground truth vector field $u(\bm{z}_{t})$ via mean squared error as follows:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{T}(0,1),\bm{z}_{t}}||v_{\theta}(\bm{z}_{t})-u(\bm{z}_{t})||^{2},$ (5)
where $\mathcal{T}(0,1)$ is a distribution ranging from 0 to 1.
However, we cannot obtain the ground truth vector field $u(\bm{z}_{t})$ directly. Ripman et al.  [36] suggest defining the conditional ground truth $u(\bm{z}_{t}|\hat{\bm{z}}_{1})$ instead. Specifically, it is modeled as a straight vector field $\hat{\bm{z}}_{1}-\bm{z}_{0}$ in Rectified Flow [13, 19]. This is called conditional flow matching [36] trained by the following objective:
@@
The gradients of $\mathcal{L}_{\text{FM}}$ of Equation 5 and $\mathcal{L}_{\text{CFM}}$ of Equation 6 are identical w.r.t $\bm{\theta}$. We exploit expressive CNFs with efficient Flow Matching training to estimate the future motion density $p(\bm{x}|\bm{c})$.
@@ 4.1 Overview of CacheFlow
We estimate the future motion density $p(\bm{x}|\bm{c})$ by transforming a conditional base distribution $q_{\phi}(\bm{z}|\bm{c})$. This $q_{\phi}$ is conditioned on the past motion $\bm{c}$. Then we can sample predictions $\bm{x}\sim p(\bm{x}|\bm{c})$ for stochastic human motion prediction. Most traditional approaches based on conditional generative models use a trivial source distribution, often a simple Gaussian. However, we redefine the source distribution to be more informative and directly regressed from past motions. This allows us to develop a much lighter and faster model for predicting future movements.
@@
This parametrization trick differs from the widely-used conditional density formulation [67] where $\bm{c}$ is conditioned to the flow-based model $f_{\theta}$. In this formulation, only the conditional base density $q(\bm{z}|\bm{c})$ varies depending on $\bm{c}$ during inference, whereas the unconditional flow-based model $\bm{x}=f_{\theta}(\bm{z})$ and the Jacobian $|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1}$ are kept same during inference and thus can be reused as-is once calculated.
Therefore, we could precompute the mapping results and Jacobians of an unconditional flow-based model $f_{\theta}$. We cache the triplets $t=\{\bm{z},|{\rm det}\mathcal{J}_{f_{\theta}}(z)|^{-1},\bm{x}\}$ for later reuse in the inference stage, as shown in Figure 2(a).
Then, during inference, we design a new trick to reuse the cached triplets by associating them with the specific conditions of the past motion sequences, as shown in Figure 2(b). Now, instead of a typical conditional generative model, e.g., conditional normalizing flow, we only need a lightweight model to model the conditional base density $q_{\phi}(\bm{z}|\bm{c})$ and achieve similar expressivity. We could finally estimate the future motion density by $p(\bm{x}|\bm{c})=q_{\phi}(\bm{z}|\bm{c})|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1}$. The method is summarized as pseudocode in Algorithm 1. In the following paragraphs, we elaborate on the details of our method.
@@ 4.2 Precompute Unconditional Flow-based Model
As the first step of our method, we use the human motion dataset to learn an unconditional flow-based model $f_{\theta}$. From this unconditional human motion prediction model, we will collect the triplets $t=\{\bm{z},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1},\bm{x}\}$ for later use. This part is illustrated in Figure 2(a).
In our implementation, we built the unconditional flow model by CNFs due to its proven expressivity for predicting human motion. The unconditional model is trained to predict a fixed-length future motion $\bm{x}$ given a noise sample $\bm{z}$ from a source distribution $q_{\phi}(\bm{z}|\bm{c})$:
@@
Because $\bm{z}$ is sampled from a known distribution and normalizing-flow models are deterministic with reversible bijective transformation, we could know the density of each $\{\bm{z},\bm{x}\}$ pair. We train the unconditional continuous normalizing flow with the flow matching objective described in Equation 6. Then we collect $K$ samples denoted by the triplet $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$. Triplets are collected by applying the inverse transform of $f_{\theta}$ to ground truth future motions in the training split. These triplets are cached for fast inference as described in Section 4.3. This caching operation is different from anchor-based methods [69, 68] since CacheFlow caches all motions of the training split.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The task of 3D human motion prediction is to forecast the future 3D pose sequence given an observed past sequence. Traditional motion prediction methods are often based on deterministic models and can struggle to capture the inherent uncertainty in human movement. Recently, stochastic approaches have addressed this limitation. Stochastic approaches allow models to sample multiple possible future motions. Stochastic human motion prediction methods utilize conditional generative models such as generative adversarial networks (GANs) (gan), variational autoencoders (VAEs) (vae), and denoising diffusion probabilistic model (ddpm). However, many stochastic approaches cannot explicitly model the probability density distribution.
Conversely, density estimate-based approaches explicitly model the probability density distribution. In safety-critical applications such as autonomous driving (survey_autonomous_driving) and human-robot interaction (koppula2015anticipating; lasota2017multiple; butepage2018anticipating), a density estimate can represent all possible future motions (not just a few samples) by tracking the volume of density. It can be used to derive guarantees on safety (mainprice2013human; sanderud2015likelihood; tisnikar2024probabilistic).
However, previous density estimation suffers from high computational cost. The expensive computational cost can prohibit applications to real-time use-cases, especially with high dimensional data such as human motions. For instance, kernel density estimation (KDE) (rosenblatt1956remarks; parzen1962estimation) requires an exponentially growing number of samples for accurate estimation. Concretely, more than one trillion samples are required for accurate KDE over a 48-dim pose over 100 frames of human motion prediction (silverman2018density).
In contrast to traditional KDE, recent parametric density estimation approaches use conditional flow-based generative models, including normalizing flows (rezende2015variational; tabak2013family; tabak2010density) and continuous normalizing flows (chen2018neural). These flow-based generative models (“flow-based model” for brevity) directly estimate the density to avoid time-consuming sampling required in KDE. However, inferring the exact probability of possible future motions remains computationally expensive. This is because capturing the full shape of the distribution requires evaluating the probabilities of many potential future motions.
To address this computational limitation, we propose a fast density estimation method based on a flow-based model called "CacheFlow". Our CacheFlow utilizes an unconditional flow-based model for prediction, as illustrated in Fig.˜1. Since the unconditional flow-based model is independent of past observed motions, its calculation can be precomputed and skipped at inference. This precomputation omits a large portion of computational cost. To achieve further acceleration, our unconditional flow-based model represents transformation between a lightweight conditional base density and the density of future motions. At inference, the density of future motion is estimated by computing the lightweight conditional base density and combining it with the precomputed results of the flow-based model. The inference of our method is approximately one millisecond.
@@
CacheFlow demonstrates comparable accuracy to previous methods on standard stochastic human motion prediction benchmarks, Human3.6M (ionescu2013h36m) and AMASS (mahmood2019amass). Furthermore, our method estimates density more accurately than previous stochastic human motion prediction methods with KDE. CacheFlow shows improved computational efficiency, making it well-suited for real-time applications. The contributions of this paper are fourfold as follows:
@@ 2.1 Human Motion Prediction
Deterministic approaches. Early approaches on human motion prediction (aksan2021spatio; bouazizi2022motionmixer; fragkiadaki2015recurrent; butepage2017deep; guo2023back; jain2016structural; li2018convolutional) focused on deterministic settings. They predict the most likely motion sequence based on the past motion. A wide range of architectures were proposed including multi-layer perceptron (guo2023back), recurrent neural networks (fragkiadaki2015recurrent; jain2016structural; martinez2017human; gui2018adversarial; pavllo2018quaternet; liu2019towards), convolutional neural networks (li2018convolutional; medjaouri2022hr), transformers (aksan2021spatio; cai2020learning; martinez2021pose), and graph neural networks (GNNs) (mao2019learning; li2020dynamic; dang2021msr; li2021skeleton). GNN can account for the explicit tree expression of the human skeleton, while other architectures implicitly learn the dependencies between joints.
Stochastic approaches. To capture the inherent uncertainty in human movements, recent works have focused on stochastic human motion prediction to predict multiple likely future motions. The main stream of stochastic methods use generative models for the purpose, such as generative adversarial networks (GANs) (barsoum2018hp; kundu2019bihmp), variational autoencoder (VAE) (walker2017pose; yan2018mt; mao2021generating; cai2021unified), and denoising diffusion probabilistic model (DDPM) (barquero2023belfusion; chen2023humanmac; wei2023human; sun2024comusion). To improve the diversity of predictions, diversity-promoting loss (mao2021generating; barquero2023belfusion) or explicit sampling techniques (wei2023human) were proposed. In contrast to generative models, anchor-based methods (anchor; xu2024learning) learn a fixed number of anchors corresponding to each prediction to ensure diversity. However, most stochastic methods cannot describe the density of future motions explicitly. This prevents exhaustive or maximum likelihood sampling for practical applications. On the contrary, our method allows for explicit density estimation using normalizing flows (kobyzev2020normalizing).
@@ 2.2 Density Estimation
Non-parametric Approach. The representative non-parametric density estimation is kernel density estimation (KDE) (rosenblatt1956remarks; parzen1962estimation). KDE can estimate density by using samples from generative models. However, KDE requires a large number of samples for accurate estimation. Therefore, it often cannot run in real-time.
Parametric Approach. As a representative parametric model, Gaussian mixture models (GMMs) parametrize density with several Gaussian distributions and their mixture weights. Its nature of mixing Gaussian priors limits its ability to generalize to complex data distribution. Another parametric approach with more expressivity is flow-based generative models (kobyzev2020normalizing). By a learned bijective process, normalizing flows (NFs) (rezende2015variational; tabak2013family; tabak2010density) transform a simple density like the standard normal distribution into a complex data density. Recently, continuous normalizing flows (CNFs) (chen2018neural; DBLP:conf/iclr/GrathwohlCBSD19) achieve more expressive density than standard normalizing flows via an ODE-based bijective process. While training of CNFs is inefficient due to the optimization of ODE solutions, an efficient training strategy named flow matching (DBLP:conf/iclr/LipmanCBNL23) was proposed. FlowChain (maeda2023fast) was proposed for fast and efficient density estimation in human trajectory forecasting. FlowChain improves the inference time efficiency by reusing results from the conditional flow-based method while the past sequences are similar. However, with significantly different past sequences, FlowChain’s efficiency can’t hold anymore. Unlike FlowChain, our method can perform fast and efficient inference regardless of past sequences.
2.3 Diffusion-based Motion Modeling
Recently, diffusion-based generative models have achieved state-of-the-art performance in synthesizing diverse and high-fidelity human motion. The Motion Diffusion Model (MDM) (mdm) demonstrated that diffusion processes can model complex kinematic distinctness, capable of tasks ranging from text-to-motion to motion completion. Concurrently, approaches inspired by Large Language Models, such as MotionGPT (motiongpt) and MotionLM (motionlm), treat motion sequences as discrete tokens, leveraging the transformer architecture to capture long-term semantic dependencies.
While these methods excel in generation quality and semantic control, they often incur high computational costs during inference. Diffusion models typically require multi-step iterative denoising, and LLM-based approaches rely on heavy autoregressive sampling. In contrast to these heavy generative frameworks, our work focuses specifically on the efficiency-latency trade-off. While we share the goal of modeling complex motion distributions, CacheFlow leverages the tractable likelihoods of Normalizing Flows and introduces caching mechanisms to ensure rapid inference suitable for real-time prediction scenarios, avoiding the computational bottleneck of iterative diffusion or large-scale transformers.
@@ 3.2 Normalizing Flow
Normalizing flow (rezende2015variational; tabak2013family; tabak2010density) is a generative model with explicit density estimation. It follows a bijective mapping $f_{\theta}$ with learnable parameters $\theta$. It transforms a simple base density $q(\bm{z})$ such as a Gaussian distribution into the complex data density $p(\bm{x})$. We can analytically estimate the exact probability via the change-of-variables formula as follows:
@@
[…]
@@ 3.3 Continuous Normalizing Flow (CNF)
Continuous normalizing flow (CNF) (chen2018neural; DBLP:conf/iclr/GrathwohlCBSD19) is a normalizing flow variant based on an ordinary differential equation (ODE). CNF defines $t$-continuous path $\bm{z}_{t}$ between the base density space $\bm{z}_{0}\sim q(\bm{z})$ and the data space $\bm{z}_{1}=\bm{x}\sim p(\bm{x})$. This $\bm{z}_{t}$ is defined by the parameterized vector field $\frac{d\bm{z}_{t}}{dt}=v_{\theta}(\bm{z}_{t})$. The data $\bm{x}=\bm{z}_{1}$ is generated via numerical integration of vector field $v_{\theta}(\bm{z}_{t})$ as follows:
$\bm{x}=\bm{z}_{1}=\bm{z}_{0}+\int_{0}^{1}v_{\theta}(\bm{z}_{t})dt.$ (4)
The CNF transformation Eq.˜4 is denoted as $\bm{x}=f_{\theta}(\bm{z})$ for brevity. Although CNF can be trained by minimizing negative log-likelihood, it is time-consuming due to the numerical integration of ODE.
@@ 3.4 Flow Matching
In order to train the parameterized vector field efficiently, one can leverage the flow matching (DBLP:conf/iclr/LipmanCBNL23) strategy. As a new training strategy, Flow Matching avoids the numerical integration of ODE by directly optimizing the vector field $v_{\theta}(\bm{z}_{t})$. The objective of flow matching is to match the parameterized vector field $v_{\theta}(\bm{z}_{t})$ to the ground truth vector field $u(\bm{z}_{t})$ via mean squared error as follows:
[…]
where $\mathcal{T}(0,1)$ is a distribution ranging from 0 to 1, such as a uniform distribution.
However, we cannot obtain the ground truth vector field $u(\bm{z}_{t})$ directly. DBLP:conf/iclr/LipmanCBNL23 suggest defining the conditional ground truth $u(\bm{z}_{t}|\hat{\bm{z}}_{1})$ instead. Specifically, it is modeled as a straight vector field $\hat{\bm{z}}_{1}-\bm{z}_{0}$ in Rectified Flow (DBLP:conf/iclr/LiuG023). This is trained by the following objective:
@@
The gradients of $\mathcal{L}_{\text{FM}}$ of Eq.˜5 and $\mathcal{L}_{\text{CFM}}$ of Eq.˜6 are identical w.r.t $\bm{\theta}$. We exploit expressive CNFs with efficient Flow Matching training to estimate the future motion density $p(\bm{x}|\bm{c})$.
@@ 4.1 Overview of CacheFlow
As mentioned in Sec. 3.2, we propose a flow-based method on a latent representation using frozen $\mathcal{E}$ and $\mathcal{D}$. We estimate the future motion density $p(\bm{x}|\bm{c})$ by transforming a conditional base distribution $q_{\phi}(\bm{z}|\bm{c})$, . This $q_{\phi}$ is conditioned on the past motion $\bm{c}$. Then we can sample predictions $\bm{x}\sim p(\bm{x}|\bm{c})$ for stochastic human motion prediction. Most traditional approaches based on conditional generative models use a trivial source distribution, often a simple Gaussian. However, we redefine the source distribution to be more informative and directly regressed from past motions. This allows us to develop a much lighter and faster model for predicting future movements.
@@
This parametrization trick differs from the widely-used conditional density formulation (conditional_normalizing_flow) where $\bm{c}$ is conditioned to the flow-based model $f_{\theta}$. In this formulation, only the conditional base density $q(\bm{z}|\bm{c})$ varies depending on $\bm{c}$ during inference, whereas the unconditional flow-based model $\bm{x}=f_{\theta}(\bm{z})$ and the Jacobian $|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1}$ are kept same during inference and thus can be reused as-is once calculated.
Therefore, we could precompute the mapping results and Jacobians of an unconditional flow-based model $f_{\theta}$. We cache the triplets $t=\{\bm{z},|{\rm det}\mathcal{J}_{f_{\theta}}(z)|^{-1},\bm{x}\}$ for later reuse in the inference stage, as shown in Fig.˜2(a).
Then, during inference, we design a new trick to reuse the cached triplets by associating them with the specific conditions of the past motion sequences, as shown in Fig.˜2(b). Now, instead of a typical conditional generative model, e.g., conditional normalizing flow, we only need a lightweight model to model the conditional base density $q_{\phi}(\bm{z}|\bm{c})$ and achieve similar expressivity. We could finally estimate the future motion density by $p(\bm{x}|\bm{c})=q_{\phi}(\bm{z}|\bm{c})|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1}$. The method is summarized as pseudocode in Algorithm˜1. In the following paragraphs, we elaborate on the details of our method.
@@ 4.2 Precompute Unconditional Flow-based Model
As the first step of our method, we use the human motion dataset to learn an unconditional flow-based model $f_{\theta}$. From this unconditional human motion prediction model, we will collect the triplets $t=\{\bm{z},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z})|^{-1},\bm{x}\}$ for later use. This part is illustrated in Fig.˜2(a).
In our implementation, we built the unconditional flow model by CNFs due to its proven expressivity for predicting human motion. The unconditional model is trained to predict a fixed-length future motion $\bm{x}$ given a noise sample $\bm{z}$ sampled from a source distribution $q(\bm{z})$, a standard Gaussian distribution:
@@
Because $\bm{z}$ is sampled from a known distribution and normalizing-flow models are deterministic with reversible bijective transformation, we could know the density of each $\{\bm{z},\bm{x}\}$ pair. We train the unconditional continuous normalizing flow with the flow matching objective described in Eq.˜6. Then we collect $K$ samples denoted by the triplet $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$. Triplets are collected by applying the inverse transform of $f_{\theta}$ to ground truth future motions in the training split. These triplets are cached for fast inference as described in Sec.˜4.3. This caching operation is different from anchor-based methods (anchor; xu2024learning) since CacheFlow caches all motions of the training split.